# Perturbation-based Learning for Recurrent Neural Networks

### v1

@@ Abstract
Recurrent neural networks (RNNs) hold immense potential for computations due to their Turing completeness and sequential processing capabilities, yet existing methods for their training encounter efficiency challenges. Backpropagation through time (BPTT), the prevailing method, extends the backpropagation (BP) algorithm by unrolling the RNN over time. However, this approach suffers from significant drawbacks, including the need to interleave forward and backward phases and store exact gradient information. Furthermore, BPTT has been shown to struggle with propagating gradient information for long sequences, leading to vanishing gradients. An alternative strategy to using gradient-based methods like BPTT involves stochastically approximating gradients through perturbation-based methods. This learning approach is exceptionally simple, necessitating only forward passes in the network and a global reinforcement signal as feedback. Despite its simplicity, the random nature of its updates typically leads to inefficient optimization, limiting its effectiveness in training neural networks. In this study, we present a new approach to perturbation-based learning in RNNs whose performance is competitive with BPTT, while maintaining the inherent advantages over gradient-based learning. To this end, we extend the recently introduced activity-based node perturbation (ANP) method to operate in the time domain, leading to more efficient learning and generalization. Subsequently, we conduct a range of experiments to validate our approach. Our results show similar performance, convergence time and scalability when compared to BPTT, strongly outperforming standard node perturbation and weight perturbation methods. These findings suggest that perturbation-based learning methods offer a versatile alternative to gradient-based methods for training RNNs.
@@ 1 Introduction
Recurrent neural networks (RNNs), with their ability to process sequential data and capture temporal dependencies, have found applications in tasks such as natural language processing (Cho et al.,, 2014; Yao et al.,, 2013; Sutskever et al.,, 2014) and time series prediction (Hewamalage et al.,, 2021). They hold immense potential for computation thanks to their Turing completeness (Chung and Siegelmann,, 2021). Furthermore, due to their sequential processing capabilities, they offer high versatility to process variable-sequence length inputs and fast inference on long sequences (Orvieto et al.,, 2023). Nevertheless, traditional training methods like backpropagation through time (BPTT) are challenging to apply (Bengio et al.,, 1994; Lillicrap and Santoro,, 2019), particularly with long sequences. Unrolling the RNN over time for gradient propagation and weight updating proves computationally demanding and difficult to parallelize with variable-length sequences. Additionally, employing BPTT can result in issues like vanishing or exploding gradients (Pascanu et al.,, 2013). Moreover, the non-locality of their updates can pose challenges when implemented on unconventional computing platforms (Kaspar et al.,, 2021).
@@
An alternative approach to training neural networks is stochastically approximating gradients through perturbation-based methods (Spall,, 1992; Widrow and Lehr,, 1990; Werfel et al.,, 2003). In this type of learning, synaptic weights are adjusted based on the impact of introducing perturbations into the network. When the perturbation enhances performance, the weights are strengthened, driven by a global reinforcement signal, and vice versa. This method is computationally simple, relying solely on forward passes and a reinforcement signal distribution across the network. It differs from gradient-based methods, such as BP, which needs a specific feedback circuit to propagate specific signals and a dedicated backward pass to compute explicit errors. This simplicity is especially beneficial for RNNs as it eliminates the need to unroll the network over time during training (Werbos,, 1990). Examples of gradient-based and perturbation-based approaches to update the neural weights are visually depicted in Figure 1. Standard perturbation methods include node perturbation (NP), where the perturbations are added into the neurons, and weight perturbation (WP), where the perturbations are added into the synaptic weights (Werfel et al.,, 2003; Züge et al.,, 2023).
@@
Nevertheless, the stochastic nature of perturbation-based updates can lead to inefficiencies in optimizing the loss landscape (Lillicrap et al.,, 2020), resulting in prolonged convergence times (Werfel et al.,, 2003). Additionally, perturbation-based often exhibit poor scalability, leading to inferior performance compared to gradient-based methods (Hiratani et al.,, 2022). This performance gap increases as the network size increases, and in large networks, instability often arises manifesting as extreme weight growth (Hiratani et al.,, 2022). Due to these challenges, a number of recent implementations have aimed to establish perturbation-based learning as an effective gradient-free method for neural network training. In (Lansdell et al.,, 2019), NP is utilized to approximate the feedback system in feedforward and convolutional networks, mitigating the error computation aspect of the BP algorithm. However, a dedicated feedback system is still necessary to propagate synapse-specific errors. In (Züge et al.,, 2023), standard NP and WP are employed on temporally extended tasks in RNNs, though direct comparisons with gradient-based algorithms like BP are lacking. Their results suggests that WP may outperform NP in specific cases.
@@
Figure 1: Gradient-based vs perturbation-based learning. Example depicts networks unrolled across 3 time steps. a) General procedure followed by gradient-based learning approaches. Sequential computation of the forward and backward passes is necessary to calculate updates. b) General procedure utilized by perturbation-based learning approaches. The computation of the eligibility trace varies based on the employed algorithm (e.g., NP, WP, ANP). In perturbation-based learning, the forward pass and noisy forward pass can be parallelized when employing two models.
In this study, we present an implementation of perturbation-based learning in RNNs whose performance is competitive with BP, while maintaining the inherent advantages over gradient-based learning. To this end, we extend the decorrelated activity-based node perturbation (DANP) approach (Dalm et al.,, 2023) to operate in the time domain using RNNs. This approach relies solely on neural activities, eliminating the need for direct access to the noise process, and incorporates a mechanism to decorrelate layer-wise inputs. The resulting updates align more closely with directional derivatives, compared to standard NP, approximating SGD more accurately. As a result, this approach significantly outperforms the standard node perturbation in practical tasks. Furthermore, input decorrelation has been shown to significantly speed up training of deep neural networks (Dalm et al.,, 2024) as uncorrelated data variables align SGD updates with the natural gradient (Desjardins et al.,, 2015). Additionally, we extend standard implementations of WP and NP to operate in the time domain, using them as baselines along with BPTT, referred to as BP in this study for simplicity.
@@ 2.1 Recurrent neural network model
The output error at each time step is defined as $\ell_{t}=||y_{t}-y^{*}_{t}||^{2}$ where $y_{t}^{*}$ denotes the target output at time $t$. The loss associated with output $y=(y_{1},\ldots,y_{T})$ with $T$ the time horizon is defined as the mean of the output errors over the entire sequence:
@@ 2.2 Node perturbation through time
where $\tilde{x}_{t}$ and $\tilde{y}_{t}$ denote the noisy neural outputs at time $t$. The noise added in the input and hidden layers is generated from zero-mean, uncorrelated Gaussian random variables $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{x})$ and $\nu_{t}\sim\mathcal{N}(0,\sigma^{2}I_{y})$, respectively, where $I_{x}$ and $I_{y}$ are identity matrices with the dimensions of the hidden and output layers. The noise injected is different in every timestep $t$.
[…]
$\Delta A$ $=\sigma^{-2}\delta\mathcal{L}\sum_{t=1}^{T}\xi_{t}x^{\top}_{t}\,,\quad\Delta B=\sigma^{-2}\delta\mathcal{L}\sum_{t=1}^{T}\nu_{t}u^{\top}_{t}\,,\quad\Delta R=\sigma^{-2}\delta\mathcal{L}\sum_{t=1}^{T}\xi_{t}x^{\top}_{t-1}$ (1)
[…]
$\Delta A$ $=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\xi_{t}x^{\top}_{t}\,,\quad\Delta B=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\nu_{t}u^{\top}_{t}\,,\quad\Delta R=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\xi_{t}x^{\top}_{t-1}$ (2)
Here, both the clean standard pass and the noisy pass could run concurrently using two identical copies of the model, which enables compatibility with online learning setups. This technique allows our method to compute and implement updates online at every time step, reducing memory requirements as updates are applied immediately without the need for storage.
Zenke and Neftci, (2020) investigate setups similar to the one proposed here, seeking to bridge the real-time recurrent learning (RTRL) algorithm (Williams and Zipser,, 1989), which is more effective for online setups than BPTT but computationally demanding, with biologically plausible learning rules. They demonstrate that by combining learning algorithms that approximate RTRL with temporally local losses, effective approximations can be achieved. These approximations notably decrease RTRL’s computational cost while preserving strong learning performance. A similar rationale is used by Bellec et al., (2020) in the context of recurrent spiking neural networks. Appendix A provides a comparison between the conventional implementation and our implementation of NP through time.
@@ 2.3 Activity-based node perturbation through time
Activity-based node perturbation (ANP) is a variant of the node perturbation approach, proposed by Dalm et al., (2023), which has been exclusively applied in feedforward networks. This approach approximates the directional derivatives across the network, resulting in a closer alignment between the updates generated by this method and those provided by BP. Additionally, it does not require direct access to the noise process itself as it operates solely by measuring changes in neural activity. Given that NP can be interpreted as a noisy variant of SGD (Hiratani et al.,, 2022), ANP can be seen as a more precise approximation of SGD. For a detailed derivation, we refer to (Dalm et al.,, 2023).
Similar to the node perturbation approach, the noisy pass is the same as the one performed in the standard node perturbation approach. Consistent with our node perturbation implementation extended over time, we calculate reinforcement signals at each time step to drive synaptic changes. Let $\alpha_{t}=Au_{t}+Rx_{t-1}$ and $\beta_{t}=Bx_{t}$ are the pre-activations in the clean pass. Similarly, let $\tilde{\alpha}_{t}=Au_{t}+R\tilde{x}_{t-1}+\xi_{t}$ and $\tilde{\beta}_{t}=B\tilde{x}_{t}+\nu_{t}$ denote the pre-activations in the noisy pass. Define $N$ as the total number of neurons in the network. We compute the learning signals responsible for weight updating as
[…]
@@ 2.4 Weight perturbation through time
As in node perturbation in the time domain, typical implementations of weight perturbation in the time domain, such as  (Züge et al.,, 2023; Cauwenberghs,, 1992), the reinforcement signal is derived from the difference in loss $\delta\mathcal{L}=\tilde{\mathcal{L}}-\mathcal{L}$, computed over the entire sequence. According to this method, the learning signals for updating the weights are defined as
$\Delta A$ $=\sigma^{-2}\delta\mathcal{L}\xi_{t}\,,\quad\Delta B=\sigma^{-2}\delta\mathcal{L}\zeta_{t}\,,\quad\Delta R=\sigma^{-2}\delta\mathcal{L}\nu_{t}$ (4)
This approach inherits the same set of drawbacks and benefits as seen in node perturbation. Hence, we again employ reinforcement signals, $\delta\ell_{t}$, computed at each time step, to drive synaptic changes. We define the learning signals to update the weights in the time domain as
$\Delta A$ $=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\xi_{t}\,,\quad\Delta B=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\zeta_{t}\,,\quad\Delta R=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\nu_{t}$ (5)

### v3

# Gradient-Free Training of Recurrent Neural Networks using Random Perturbations

@@ Abstract
Recurrent neural networks (RNNs) hold immense potential for computations due to their Turing completeness and sequential processing capabilities, yet existing methods for their training encounter efficiency challenges. Backpropagation through time (BPTT), the prevailing method, extends the backpropagation (BP) algorithm by unrolling the RNN over time. However, this approach suffers from significant drawbacks, including the need to interleave forward and backward phases and store exact gradient information. Furthermore, BPTT has been shown to struggle to propagate gradient information for long sequences, leading to vanishing gradients. An alternative strategy to using gradient-based methods like BPTT involves stochastically approximating gradients through perturbation-based methods. This learning approach is exceptionally simple, necessitating only forward passes in the network and a global reinforcement signal as feedback. Despite its simplicity, the random nature of its updates typically leads to inefficient optimization, limiting its effectiveness in training neural networks. In this study, we present a new approach to perturbation-based learning in RNNs whose performance is competitive with BPTT, while maintaining the inherent advantages over gradient-based learning. To this end, we extend the recently introduced activity-based node perturbation (ANP) method to operate in the time domain, leading to more efficient learning and generalization. We subsequently conduct a range of experiments to validate our approach. Our results show similar performance, convergence time and scalability when compared to BPTT, strongly outperforming standard node perturbation and weight perturbation methods. These findings suggest that perturbation-based learning methods offer a versatile alternative to gradient-based methods for training RNNs which can be ideally suited for neuromorphic computing applications.
@@ 1 Introduction
Recurrent neural networks (RNNs), with their ability to process sequential data and capture temporal dependencies, have found applications in tasks such as natural language processing (Cho et al.,, 2014; Yao et al.,, 2013; Sutskever et al.,, 2014) and time series prediction (Hewamalage et al.,, 2021). They hold immense potential for computation thanks to their Turing completeness (Chung and Siegelmann,, 2021). Furthermore, due to their sequential processing capabilities, they offer high versatility to process variable-sequence length inputs and fast inference on long sequences (Orvieto et al.,, 2023). Nevertheless, traditional training methods like backpropagation through time (BPTT) (Werbos,, 1990) are challenging to apply (Bengio et al.,, 1994; Lillicrap and Santoro,, 2019), particularly with long sequences. Unrolling the RNN over time for gradient propagation and weight updating proves to be computationally demanding and difficult to parallelize with variable-length sequences. Additionally, employing BPTT can result in issues like vanishing or exploding gradients (Pascanu et al.,, 2013). Moreover, the non-locality of their updates can pose significant challenges when implemented on unconventional computing platforms (Kaspar et al.,, 2021).
@@
An alternative approach to training neural networks is stochastically approximating gradients through perturbation-based methods (Spall,, 1992; Widrow and Lehr,, 1990; Werfel et al.,, 2003). In this type of learning, synaptic weights are adjusted based on the impact of introducing perturbations to the network. When the perturbation enhances performance, the weights are strengthened, driven by a global reinforcement signal, and vice versa. This method is computationally simple, relying solely on forward passes and a reinforcement signal distribution across the network. It differs from gradient-based methods, such as BP, which needs a specific feedback circuit to propagate specific signals and a dedicated backward pass to compute explicit errors. This simplicity is especially beneficial for RNNs, as it eliminates the need to unroll the network over time during training (Werbos,, 1990). Examples of gradient-based and perturbation-based approaches to update the neural weights are visually depicted in Figure 1. Standard perturbation methods include node perturbation (NP), where the perturbations are added into the neurons, and weight perturbation (WP), where the perturbations are added into the synaptic weights (Werfel et al.,, 2003; Züge et al.,, 2023).
@@
Nevertheless, the stochastic nature of perturbation-based updates can lead to inefficiencies in optimizing the loss landscape (Lillicrap et al.,, 2020), resulting in prolonged convergence times (Werfel et al.,, 2003). Additionally, perturbation-based often exhibit poor scalability, leading to inferior performance when compared to gradient-based methods (Hiratani et al.,, 2022). This performance gap increases as the network size increases, and in large networks, instability often arises manifesting as extreme weight growth (Hiratani et al.,, 2022). Due to these challenges, a number of recent implementations have aimed to establish perturbation-based learning as an effective gradient-free method for neural network training. In (Lansdell et al.,, 2019), NP is utilized to approximate the feedback system in feedforward and convolutional networks, mitigating the error computation aspect of the BP algorithm. However, a dedicated feedback system is still necessary to propagate synapse-specific errors. In (Züge et al.,, 2023), standard NP and WP are employed on temporally extended tasks in RNNs, though direct comparisons with gradient-based algorithms like BP are lacking. Their results suggest that WP may outperform NP in specific cases.
@@
Figure 1: Gradient-based vs perturbation-based learning. Example depicts networks unrolled across 3 time steps. a) General procedure followed by gradient-based learning approaches. Sequential computation of the forward and backward passes is necessary to calculate updates. b) General procedure utilized by perturbation-based learning approaches. The computation of the eligibility trace varies based on the employed algorithm (e.g., NP, WP, ANP). In perturbation-based learning, the forward pass and noisy forward pass can be parallelized by employing two models.
In this study, we present an implementation of perturbation-based learning in RNNs whose performance is competitive with BP, while maintaining the inherent advantages over gradient-based learning. To this end, we extend the activity-based node perturbation (ANP) approach (Dalm et al.,, 2023) to operate in the time domain using RNNs. This approach relies solely on neural activities, eliminating the need for direct access to the noise process. The resulting updates align more closely with directional derivatives, compared to standard NP, approximating SGD more accurately. As a result, this approach significantly outperforms the standard NP in practical tasks. Furthermore, we extend standard implementations of NP and WP to operate in the time domain, using them as baselines along with BPTT, referred to as BP in this study for simplicity. In addition, we augment our time-extended implementations of ANP, NP, and WP (as well as BP) with a decorrelation mechanism, as done in the original ANP work (Dalm et al.,, 2023). This decorrelation mechanism has been shown to significantly accelerate the training of deep neural networks (Dalm et al.,, 2024) by aligning SGD updates with the natural gradient through the use of uncorrelated input variables (Desjardins et al.,, 2015). We name the resulting decorrelated methods DANP, DNP, DWP and DBP.
@@ 2 Methods
In this section, we describe our original time-extended implementations of existing perturbation-based algorithms for feedforward networks, adapted for compatibility with RNNs and time-extended tasks. We highlight how our implementations differ from other time-extended methods in the literature. For clarity and ease of understanding, we present them in the following order: Node perturbation (NP), activity-based node perturbation (ANP), and weight perturbation (WP). These three methods will throughout be compared to backpropagation (BP). BP, WP and NP will throughout be used as baseline methods, with ANP (extended to the time-domain) being our novel contribution. Additionally, we describe the decorrelation mechanism proposed by Dalm et al., (2023) for feedforward networks, which we incorporate into our RNN model to enhance the learning efficiency of the algorithms under consideration.
@@ 2.1 Recurrent neural network model
The output error at each time step is defined as $\ell_{t}=||y_{t}-y^{*}_{t}||^{2}$ where $y_{t}^{*}$ denotes the target output at time $t$. The loss associated with output $y=(y_{1},\ldots,y_{T})$, with $T$ the time horizon, is defined as the mean of the output errors over the entire sequence:
@@ 2.2 Node perturbation through time
where $\tilde{x}_{t}$ and $\tilde{y}_{t}$ denote the noisy neural outputs at time $t$. The noise added in the hidden and output layers is generated from zero-mean, uncorrelated Gaussian random variables $\xi_{t}\sim\mathcal{N}(0,\sigma^{2}I_{x})$ and $\nu_{t}\sim\mathcal{N}(0,\sigma^{2}I_{y})$, respectively, where $I_{x}$ and $I_{y}$ are identity matrices with the dimensions of the hidden and output layers. The noise injected is different in every timestep $t$.
In typical implementations of node perturbation in the time domain, as seen in works like (Züge et al.,, 2023; Fiete and Seung,, 2006), the reinforcement signal is derived from the difference in loss, $\mathcal{L}=\tilde{\mathcal{L}}-\mathcal{L}$, where $\tilde{\mathcal{L}}$ and $\mathcal{L}$ represent the loss of the noisy pass and the clean pass, respectively. This signal captures the network’s overall performance across the entire sequence. Additionally, an eligibility trace, computed as the sum over time of the pre-synaptic neuron’s output multiplied by the injected perturbation ($\xi_{t}x_{t}^{\top}$, $\nu_{t}x_{t}^{\top}$ or $\xi_{t}x_{t-1}^{\top}$, depending on the synaptic weights being updated), is utilized. According to this method, the learning signals for updating the weights are defined in Equation 1.
$\Delta A$ $=\sigma^{-2}\delta\mathcal{L}\sum_{t=1}^{T}\xi_{t}x^{\top}_{t},\quad\Delta B=\sigma^{-2}\delta\mathcal{L}\sum_{t=1}^{T}\nu_{t}u^{\top}_{t},\quad\Delta R=\sigma^{-2}\delta\mathcal{L}\sum_{t=1}^{T}\xi_{t}x^{\top}_{t-1}.$ (1)
While this approach offers benefits such as enhanced compatibility with delayed rewards, we here consider a local approach, where we only consider the local loss difference $\delta\ell_{t}=\tilde{\ell}_{t}-\ell_{t}$ in individual timesteps when computing updates in contrast to using the total loss difference $\mathcal{L}$. This reward-per-time step is employed alongside eligibility traces local on time defined as the product of the pre-synaptic neuron’s output and the injected perturbation at each time step ($\xi_{t}x_{t}^{\top}$, $\nu_{t}x_{t}^{\top}$ or $\xi_{t}x_{t-1}^{\top}$, depending on the synaptic weights being updated). According to our approach, the learning signals to update the weights over time are defined in Equation 2.
$\Delta A$ $=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\xi_{t}x^{\top}_{t},\quad\Delta B=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\nu_{t}u^{\top}_{t},\quad\Delta R=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\xi_{t}x^{\top}_{t-1}.$ (2)
Here, both the clean standard pass and the noisy pass could run concurrently using two identical copies of the model, which enables compatibility with online learning setups. This technique allows our method to compute and implement updates online at every time step.
Zenke and Neftci, (2020) investigate setups similar to the one proposed here, seeking to bridge the real-time recurrent learning (RTRL) algorithm (Williams and Zipser,, 1989), which is more effective for online setups than BPTT but computationally demanding, with biologically plausible learning rules. They demonstrate that by combining learning algorithms that approximate RTRL with temporally local losses, effective approximations can be achieved. These approximations notably decrease RTRL’s computational cost while preserving strong learning performance. A similar rationale is used by Bellec et al., (2020) in the context of recurrent spiking neural networks. We provide a comparison between the conventional implementation and our implementation of NP through time in Appendix A.
@@ 2.3 Activity-based node perturbation through time
Activity-based node perturbation (ANP) is a variant of the node perturbation approach, proposed by Dalm et al., (2023), which has been exclusively applied in feedforward networks. This approach approximates the directional derivatives across the network, resulting in a closer alignment between the updates generated by this method and those provided by BP (also see (Baydin et al.,, 2022)). Additionally, it does not require direct access to the noise process itself as it operates solely by measuring changes in neural activity. Given that NP can be interpreted as a noisy variant of SGD (Hiratani et al.,, 2022), ANP can be seen as a more precise approximation of SGD. For a detailed derivation of the link between ANP and SGD, we refer to (Dalm et al.,, 2023).
Similar to the node perturbation approach, the noisy pass is the same as the one performed in the standard node perturbation approach. Consistent with our node perturbation implementation extended over time, we calculate reinforcement signals at each time step to drive synaptic changes. Let $\alpha_{t}=Au_{t}+Rx_{t-1}$ and $\beta_{t}=Bx_{t}$ are the pre-activations in the clean pass. Similarly, let $\tilde{\alpha}_{t}=Au_{t}+R\tilde{x}_{t-1}+\xi_{t}$ and $\tilde{\beta}_{t}=B\tilde{x}_{t}+\nu_{t}$ denote the pre-activations in the noisy pass. Define $N$ as the total number of neurons in the network. We compute the learning signals responsible for weight updating as defined in Equation 3.
$\Delta A$ $=N\sum_{t=1}^{T}\ell_{t}\frac{\delta\alpha_{t}}{||\delta\alpha_{t}||^{2}}x^{\top}_{t},\quad\Delta B=N\sum_{t=1}^{T}\ell_{t}\frac{\delta\beta_{t}}{||\delta\beta_{t}||^{2}}u^{\top}_{t},\quad\Delta R=N\sum_{t=1}^{T}\ell_{t}\frac{\delta\alpha_{t}}{||\delta\alpha_{t}||^{2}}x^{\top}_{t-1},$ (3)
where $\delta\alpha_{t}=\tilde{\alpha}_{t}-\alpha_{t}$ and $\delta\beta_{t}=\tilde{\beta}_{t}-\beta_{t}$ are the pre-activation differences between the forward passes.
@@ 2.4 Weight perturbation through time
Similar to node perturbation in the time domain, typical implementations of weight perturbation in the time domain, such as those described by (Züge et al.,, 2023; Cauwenberghs,, 1992), derive the reinforcement signal from the difference in loss $\delta\mathcal{L}=\tilde{\mathcal{L}}-\mathcal{L}$, computed over the entire sequence. According to this method, the learning signals for updating the weights are defined as in Equation 4.
$\Delta A$ $=\sigma^{-2}\delta\mathcal{L}\xi_{t},\quad\Delta B=\sigma^{-2}\delta\mathcal{L}\zeta_{t},\quad\Delta R=\sigma^{-2}\delta\mathcal{L}\nu_{t}.$ (4)
This approach inherits the same set of drawbacks and benefits as seen in node perturbation. Hence, we again employ reinforcement signals, $\delta\ell_{t}$, computed at each time step, to drive synaptic changes. We define the learning signals to update the weights in the time domain as in Equation 5.
$\Delta A$ $=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\xi_{t}\,,\quad\Delta B=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\zeta_{t}\,,\quad\Delta R=\sigma^{-2}\sum_{t=1}^{T}\delta\ell_{t}\nu_{t}\,.$ (5)